August 13th 2021

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCal
      • HCal fully cabled
      • starting reestablishing DAQ, can log to ROC
      • will work today on HV (rpi20 and 21)
      • and try to take data
      • summing trigger should be working, need to adjust for delays with different PMTs
    • Bigbite Shower
      • Eric, Provokar
      • need plug in 15 more bundles
      • need plug scalers ( NIM/ECL converter to scaler and F1 TDC)
      • set up cosmics
      • need cable for BigBite shower thresholds
      • will finish cabling this week, start turning on next week
      • check mapping
      • Bryan configuring all detectors today - think to finish today
    • Bigbite Hodoscope
      • all ADC and TDC cabled at detector frame and signal going to bunker, all cables reach in bunker - TDC cables a bit short to reach converter -
      • connecting ADC, TDC, HV and LV left
      • software progress by Rachel
      • ADC cables 100 m
      • need to setup gate for ADC
    • GRINCH
      • existing cabling was redone at Bogdan's request - slack near BB was moved in near main bunker
      • half fully cable
      • other
      • TDC will be redone
      • HV cables laid and are connected
      • rpi17 HV
      • share same LV cable
      • software ok, still need online monitoring
      • still need to move to a-onl account
      • scalers display ? Existing and will be installed
    • LHRS
      • Maria learning PID for LHRS software
      • Bob helping Ole in counting room ( should be ready in 2 weeks )
      • still connect to connect to TS - should receive the part fiber next week and can try
      • pulser to simulation coincidence
      • EDTM ? Need to look for it
    • CDet
      • taking data to optimize resistors
      • after look different HV for efficiency
      • helping in Hall for installation
      • CNU starting again next week
    • GEM BB Hall A
      • 23 MPDs in Hall
      • all up to date for firmware
      • can update firmware in test lab
      • 10:30 AM power outage, because UV layers, LV was plug on same socket and tripped. Nothing can be read out yet. Being investigated, Thir bringing spares
      • need trigger fan/out box for rest of MPDs
    • GEM INFN Test Lab
      • HV test on two SBS layers, HV tripping is due to shielding - putting shielding in pieces to see where it can short -
      • SSP crate to update firmware for MPDs
    • RF time
    • HV
      • MPOD HV GUI for the GEM people
      • HV for shower and GRINCH up
      • will check hcal
    • Scalers
      • Also I got the DSC modules to work as scalers and I am setting up a bunch of SIS scalers as 3800s. I also have 4 G0 scalers that we could use, but we should make sure that they have wide input signals (>10 ns) and that we don’t use them for anything critical like deadtime, efficiency, beam current etc.
      • Bigbite scalers to be added to this crate
    • Recoil polarimeter
      • hodoscope plugged
      • DAQ rack for experiment set up
      • starting using Hall A system to boot CPU
    • EPICS
      • David found alh not available ( used on old computer - hacweb7 ) will move to adaqsc -
    • Helicity
      • scalers installed, Paul back from vacation

routine pick -up scaler event for the helicity seed and will be inserted in data stream

      • helicity signal in counting room - not reaching HRS -need to check cable and trace
      • same data stream in SBS scaler crate : need libraries on sbsvme29, library for 3801 - could be easier if in different crate - will check to put it in TS crate or hodo crate
      • reported helicity and T settle signal tagged in trigger
      • taking the RHRS signal to SBS weldement
    • start / stop / EPICS script
      • Bob has templates working with CODA3


  • On-going tasks
    • Network setup ( done ) - SBS MPD side might need few cables - need pull network for SBS side (0%)
    • Scaler crate setup (70 % - Steve ) - Steve could not talk to CPU - IP fixed with DHCP - need change fan tray - CAEN
    • documentation of VME crate names to power cycle - need to make sure all VXS are in dhcp -Bryan
    • 3 rd GEM VXS crate install router moved up ( 100% )
    • need power and network in GEM bunker ( done )
    • boot VME controllers (100 % done - need ask hostname TS crate )
    • TS cables - all fibers plugged - will post informations -hodoscope / grinch/ shower separate - can be combined at any time now everything plugged to TS
    • detector cabling
    • new MPD production - about to start - FPGA shipped to Paolo -
    • MPD VTP readout - on-going in daq lab - still having issue getting APV sending data to MPD - events going through high speed link
    • MPD being setup in Hall A - need setup VTPs to boot (done ) - 10 gigE network plugged (done) - will be able to test in Hall A - LV installed soon (done )
    • setup portservers : 1 4 port portserver for VTP (done), still need to set up a 8 port one and HCAL portserver
    • need to install remote switch for crate in bunker - rest are VXS or VME64 -
    • Bigbite - HCal trigger setup - need check with Bob if additionnal NIM crate need to be install
    • TS setup - 50 %
    • Beamline crate setup ( in scaler crate ) - FADC is in scaler need to be setup - need pull raster BPM BCM (0% will do it with David )
    • HCAL cabling
    • HCAL need power ( done, Berta )
    • Mark hodoscope trigger ?
    • Computer move to Hall A : sbsdaq1/2/3 tedbbdaq (done ) enpcamsonne ( done ) - need to install X server - Need ask Ole -
    • LHRS fiber pull ( waiting fiber August 13th ) - most likely LHRS at end of SBS run
    • Cosmics trigger ( Bogdan ? 0 %) -
    • Cosmics HCAL
    • Cosmics Shower
    • Cosmics Hodoscope
    • Test 10 gig link to DAQ to adaq1
    • EPICS script -
    • alarm handler (David)
  • Outstanding issues
    • 2 bad FADCs Bigbite - 1 FADC was repaired - one being repaired - 1 Compton FADC - repaired can use as spare ( done )
    • SD delay : still need 2 SDs ( Glasgow and INFN just requested quote )
    • need to adjust for HCAL transit time with different PMTs
    • status of cosmics paddles ( installed by Bogdan )
    • GEM fibers too short : Mark will order parts - need to check status of order
    • 15 bad MPDs - need to look at fixing - Ben ? Dustin loaning us 6 MPDs - need 85 total - 67 in hand - 1 MPD (301) VME address issue in some slot -
      • APV needed 1208 - 966 in hand working - Michael might lend 40 - ( check with Thir )
    • GEM fiber possibly bad - 2 broken - if long sleeve want to pull spare in separate conduit
    • GEM HV IP need setup ( Alex ) Done
    • GEM clock/trigger box for BB side